OpenStudy (anonymous):

PLEASE HELP: I have 38 coins. They consist of dimes and quarters. How many dimes do I have? I have $6.95

OpenStudy (amistre64):

how many dimes and quarters do you have?

OpenStudy (anonymous):

38 Total

OpenStudy (amistre64):

d + q = 38 each d is worth .10 each q is worth .25 sooo .10 d + .25 q = 6.95 you understand this?

OpenStudy (anonymous):

Yes, But it has to equal 38 coils. I got that far haha

OpenStudy (amistre64):

by substitution, let d = 19+n, and q=19-n does 19+n + 19-n = 38?

OpenStudy (anonymous):

Yes..

OpenStudy (amistre64):

then lets solve for n .10(19+n) + .25(19-n) = 6.95 when we know n we can know d and q

OpenStudy (anonymous):

95..?

OpenStudy (anonymous):

Wait.. I messed up somewhere

OpenStudy (amistre64):

we could have just as well defined it: d = 38-n, q=n .10(38-n) + .25n = 6.95 is another way to approach it ....

OpenStudy (anonymous):

21

OpenStudy (amistre64):

the idea is to set up some convenient one variable setup that defines d + q for all n the second setup, n=21 yes :)

OpenStudy (amistre64):

d = 38 - 21, q = 21

OpenStudy (amistre64):

in the first setup, n=-2 d = 19 +-2, q = 19 - -2 we get the same results

OpenStudy (anonymous):

17 dimes?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

:DDD

OpenStudy (anonymous):

I tried the same method but I can't get it right

OpenStudy (anonymous):

IT WORKED !

OpenStudy (amistre64):

:) it better worked lol

OpenStudy (amistre64):

this type of substitution is very ancient, the greeks way back when used to use it.